OpenStudy (anonymous):

The sum of two numbers is 51 and the difference is 15 . What are the numbers?

OpenStudy (trojanpoem):

assume x > y x+ y = 51 x - y = 15 2x = 66 so x = 33 y = 51 - x = 51 - 33 = 18